# Low-Cost Deterministic Networking AP access point

### v1

@@ V End Host Requirements and Measurements
The flow admission and access control mechanisms in LCDN use network calculus (NC) via Chameleon to achieve predictable latencies in the network. In the framework, it is important that any source sends, at maximum, the advertised amount of data. Mathematically, NC models the traffic sources as arrival curves. These arrival curves describe the continuous data rate $r$ and a maximum deviation $b$ at time $t$. To ensure that the sources in hosts comply with the arrival curves, LCDN uses Token-Bucket Filtering from the Traffic Control Layer in Linux’s network stack. To achieve network determinism, the output of the Token-Bucket Filter (TBF) and the parameters of the arrival curve need to be the same. The next evaluation validates the TBF’s performance in different settings. In the spirit of affordability, the test uses a Raspberry Pi 4 with 8GB of RAM. The test procedure is as follows. The Raspberry Pi connects directly to a PC with a measurement network card. Then, the TBF parameters are set on the interface. Next, the Raspberry Pi sends data to the measurement PC. Finally, the measurement PC receives the traffic and calculates the rate. Figure 4 shows the deviation of Token Bucket Parameters and the measured rate in percent. Figure 4(a) shows the deviation for different rates with a 1542 Byte burst in percent. Here, the result shows that the actual measured rate is slightly higher than the rate set in the TBF. However, the deviation is below 1.88 % and relatively constant. LCDN can model the deviation of the TBF by adjusting the rate parameter of new flows when sending it to the DNC Agent. 4(b) illustrates the deviation of the TBF with different burst sizes and a constant rate of 3 Mbit/s in percent. Here, the deviation largely depends on the burst size. Furthermore, with small burst values the deviation is large with up to 50 % deviation. With burst sizes larger than 600 Byte, the deviation is smaller than 5 %. Even with the large deviation, LCDN can model the behavior of the TBF. For this, LCDN looks at the flow request, checks the burst size, looks up the deviation in a table, changes the flow rate of the incoming request, and sends it to further processing.
Additionally, LCDN uses the VLAN-Headers for routing. Therefore, all traffic must be tagged to the correct VLAN ID and priority. Ethernet sub-interfaces allow for an easy configuration of VLAN-Headers in Linux-based systems. Therefore, LCDN’s end hosts use these sub-interfaces for VLAN tagging. Furthermore, LCDN accounts for the overhead in burst size and rate that the addition of VLAN-Headers causes.
@@
VI Discussion on Limitations
LCDN is designed to work with relatively low-cost hardware. In this initial setup, LCDN works with the FS S2805S switch series, which starts at around 85 €, and the Raspberry Pi 4, which is in a similar price range. Clearly, these devices come with limitations. These limitations, on the one side, decrease the performance of LCDN compared to larger data-center-focused systems like Chameleon [10]. On the other side, the limitations also help LCDN, especially when looking at the routing. For example, LCDN uses a single VLAN-Tag to route the packet through the network. In total, there are 4094 usable VLAN-Tags, which limit the number of flows. However, when using lower-powered end hosts, LCDN does not expect to exhaust the number of VLAN-Tags and all possible routes. Another limitation is the number of MAC-Addresses a switch can use, which is 8000 MAC-Addresses per switch.
LCDN expects that the networks and setups it uses are of reasonable size and within the capabilities of the switches and end hosts.
VII Conclusion
LCDN is a system that presents network determinism with affordable hardware. It uses a centralized controller that overviews the whole state of the network. This allows LCDN to adapt a Network Calculus framework from an already existing controller to suit its use case with low-cost devices. Using low-cost and, therefore, lower-powered devices has implications for the system, especially for the routing component. LCDN solves the routing problem with multiple spanning trees. LCDN addresses other shortcomings by having additional software components that keep track of potential violations. Then, LCDN can intervene and preserve network determinism. Furthermore, this paper provides important parameters for the NC framework for the FS S2805S switch. Additionally, this paper measures the determinism of a Raspberry Pi 4 and shows how it can be used with LCDN.
In the future, we want to compare LCDN against other systems that provide deterministic latencies and further quantify the limitations of LCDN caused by low-cost devices.

### v2

@@ VI End Host Requirements and Measurements
The following validates the TBF’s performance in different settings. In the spirit of affordability, the test uses a Raspberry Pi 4 with 8GB of RAM, assuming that any system with more computing resources and superior hardware will perform better. The test procedure is as follows: the Raspberry Pi connects directly to a PC with a measurement network card. Then, the TBF parameters are set on the interface. Next, the Raspberry Pi sends data to the measurement PC. Finally, the measurement PC receives the traffic and calculates the received rate. Figure 4 shows the deviation of Token Bucket Parameters and the measured rate in percent. Figure 4(a) shows the deviation for different rates with a 1542 Byte burst in percent. Here, the result shows that the actual measured rate is slightly higher than the rate set in the TBF, indicating that the filter is leaky. However, the deviation is below 1.88 % and relatively constant. LCDN can model the deviation of the TBF by adjusting the rate parameter of new flows when sending it to the DNC Agent. 4(b) illustrates the deviation of the TBF with different burst sizes and a constant rate of 3 Mbit/s in percent. Here, the deviation largely depends on the burst size. Furthermore, with small burst values the deviation is large with up to 50 % deviation. With burst sizes larger than 600 Byte, the deviation is smaller than 5 %. Even with the large deviation, LCDN can model the behavior of the TBF. For this, LCDN looks at the flow request, checks the burst size, looks up the deviation in a look table, changes the flow rate of the incoming request, and sends it to further processing.
@@
VII Limitations and Conclusion
LCDN is designed to work with relatively low-cost hardware. In this initial setup, LCDN works with the FS S2805S switch series, which starts at around 80€, and the Raspberry Pi 4, which is in a similar price range. Qualitatively 20€ switches also indicate that they could work with LCDN. Clearly, these devices have limitations. These limitations, decrease LCDN’s performance compared to other TSN methods. The number of usable VLANs (4094) available on the switches, slower communication with the switch’s management interface, and the overhead of configuring multiple spanning trees with constrained hardware all present challenges for using LCDN in small networks or testbeds. Moreover, the lack of advanced schedulers like TAS also restrict its use to applications where few time-sensitive flows must co-exist with other background traffic.
LCDN is a novel system that presents network determinism with affordable hardware. It uses a centralized controller that monitors the whole state of the network. This allows LCDN to adapt a Network Calculus framework from an already existing controller to suit its use case with low-cost devices. Using low-cost and, therefore, lower-powered devices has implications for the system, especially for the routing over L2 switches. LCDN solves the routing problem with multiple spanning trees and source-routed flows. LCDN addresses other integration shortcomings by having additional software components in the data plane that keep track of potential violations. That way, LCDN can intervene and preserve network determinism. This paper also provides important parameters for the NC framework for the FS S2805S switch via measurements. Additionally, this paper measures the determinism of a Raspberry Pi 4 and shows how it can be used with LCDN.
Acknowledgements
This work received funding from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 316878574.